OpenStudy (anonymous):

What is the solution to the following system? -2x-y=1 -4x-2y=-1

OpenStudy (anonymous):

Try the method that I showed you in the last one

OpenStudy (anonymous):

If you do it like the last question you will get it right!

OpenStudy (anonymous):

Can you show me the steps?

OpenStudy (anonymous):

i dont remember

OpenStudy (anonymous):

ok solve the first equation for either x or y, you pick.

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

can you solve -2x-y=1 for y?

OpenStudy (anonymous):

There's an easier way to solve this one. Use elimination.

OpenStudy (anonymous):

that is true as well

OpenStudy (anonymous):

You find the answer much quicker

OpenStudy (anonymous):

Yes, I agree :)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

you would have to either make the x's or y's equal. so lets make the x's equal. If I want both equations to have the same X values, do you know how to do that?

OpenStudy (anonymous):

She closed this one and opened it again... sO im thinking she just wants the answer instead of learning it lol

OpenStudy (anonymous):

Sucks for her. I am going to write it out anyways. Maybe she will learn two things. How to learn the stuff and how to learn to be patient. Here this is what you are going to do. You want the x's or y's cancel each other out when the equations are added together. So what you have to do is. -2x-y=1 -4x-2y=-1 Take the first equation and multiply it by -2 You get 4x + 2y = -2 Add the new equation to the second equation from the original problem. 4x + 2y = -2 + -4x - 2y = -1 ----------------- 0 0 = -3 Now you have no x and no y and it equals -3 So what do you think the answer is?

OpenStudy (anonymous):

many solutions?

OpenStudy (anonymous):

Does 0 ever = -3?

OpenStudy (anonymous):

These two equation are parallel

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it many solutions or no solutions?

OpenStudy (anonymous):

You think 0=-3? zero= -three? No never. zero is a number and -three is a number are they the same number? No not the same number and never will be.

OpenStudy (anonymous):

@Emily778 What math are you in? I believe you're in the wrong math if you didn't know that 0 does not equal -3...